Apply second substitution theorem evaluate


Integral 1 to 4 dt/(|t+4|√t)

1
Expert's answer
2022-03-29T12:04:22-0400

"I=\int_1^4\frac {dt}{|t+4|\sqrt t}"

Substitution:

"u=\sqrt t" ; "du=\frac12\frac{dt}{\sqrt t}";

"u_1=\sqrt 1=1" ; "u_2=\sqrt 4=2" ;

"I=\int_1^4\frac {dt}{|t+4|\sqrt t}=2\int_1^4\frac {dt}{2|t+4|\sqrt t}=2\int_1^2\frac {du}{|u^2+4|}=""2\int_1^2\frac {du}{u^2+4}"

Second substitution:

"v=\frac{u}{2}"; "dv=\frac{du}{2}";

"v_1=\frac12"; "v_2=1";

"I=4\int_{\frac12}^1\frac{dv}{4v^2+4}=\int_{\frac12}^1\frac{dv}{v^2+1}=\arctan{v}|_\frac12^1=""\arctan{1}-\arctan{\frac12}=\frac{\pi}{4}-\arctan{\frac12} \approx0.32"